A sample of 25.0mL of 0.105mol L-1 NaOH and a sample of 50.0mL of
0.240mol L-1 NaOH are mixed and the solution is made up to 100.00mL by adding sufficient water. What is the final concentration of this solution?
1
Expert's answer
2020-07-10T05:23:43-0400

Must work out total number of moles in total volume

In 1st sample n NaOH = cV = 0.105 X 0.025 =0.002625mol


In 2nd sample n NaOH = cV =

0.240 X 0.050 = 0.012mol


Total moles = 0.014625mol


Final concentration = n/V = 0.014625/0.100 = 0.146M